without graphing the function y = -2cos(5x) determine its amplitude, period and section width
Answer #1

Solution-

Consider the function

y = -2cos(5x) ...(1)

We know that the standard cosine function is

y = Acos(Bx) ...(2)

Here, |A| = Amplitude and Period = 2π/B

On comparing equations (1) and (2), we get

A = -2 and B = 5

So,

Amplitude = |A| = |-2| = 2

Period = 2π/B = 2π/5

Section width = Period/2 = (2π/5)/2 = π/5

Hence,

Amplitude = 2 ,Period = 2π/5 and Section Width = π/5 .